OpenStudy (anonymous):

a boy walked over hills at 3 km per hours and on flat road at 4 km per hour. He walked 18 km in 5 hours. how much of his trip was over hills?

OpenStudy (emmamink):

4 hours of his trip was over hills.

OpenStudy (emmamink):

Medal if you appreciate!!!

OpenStudy (anonymous):

how id u get that

OpenStudy (emmamink):

hold on I have to empty the dishdrainer

OpenStudy (emmamink):

So, some hours have to be at 4 miles per hour and some at three. Everything has to amount to 18 though. So, to do that we can subtract 4 from eighteen and that is one of our hours. He walked 4 miles per hour for 1 hour. Now we have 12 leftover and how many times can 3 go into that? 4 times, which is the number of hours we were looking for. So 4 hours was over the hills! I hope that wasn't too confusing.

OpenStudy (emmamink):

Well, I have to go so then if you want medal and if something doesn't make sense, tag me or message me. @EmmaMink

OpenStudy (emmamink):

hey ttb123456789

OpenStudy (anonymous):

18-4=14 not 12 3 hrs flat and 2 hrs. overhills just using common sense but how do I set up problem

jimthompson5910 (jim_thompson5910):

x = # of hours walking at a speed of 3 km/hr y = # of hours walking at a speed of 4 km/hr the boy walked a total of 5 hours, so the sum of x and y must be 5 x+y = 5 the boy walked 3km/hr for x hours, so he walked 3x km the boy walked 4km/hr for y hours, so he walked 4y km in total, he walked 3x+4y km and this total is 18 km, so 3x+4y = 18 the system of equations is x+y = 5 3x+4y = 18

OpenStudy (anonymous):

well i guess you are correect

OpenStudy (emmamink):

Oh ya oops. Sorry I screwed up. dang it

OpenStudy (emmamink):

well bye